OpenStudy (anonymous):

given f(x)=6x^2+5x-4 and g(x)=2x-1 (f/g)(x) for what value will this quotient of functions be undefined. i dont understand what she is asking for. can anyone help? @precal

OpenStudy (anonymous):

solve 2x-1=0 for x

OpenStudy (anonymous):

ok ... x=1/2 ?

OpenStudy (anonymous):

yes for x=1/2 you will get zero denominator and Quotient will be undefined.

OpenStudy (precal):

 you were given the correct solutions